# Interplay between structural, electronic and topological properties in low-dimensional tellurium

### v2

@@ 3 Results and Discussions
where $\Omega_{xy}(\mathbf{k})$ is the Berry curvature summed over the occupied bands, $\Omega_{xy}(\mathbf{k})=\sum_{n\in\mathrm{occ}}\Omega_{n,xy}(\mathbf{k})$. In our calculations, $\Omega_{n,xy}(\mathbf{k})$ is evaluated using the Kubo (velocity-matrix) expression, as implemented in WannierTools,
@@
The Berry curvature and Chern number were computed using tight-binding Hamiltonians constructed from maximally localized Wannier functions, ensuring an accurate interpolation of the electronic structure on dense $\mathbf{k}$-point meshes. Spin–orbit coupling was included in all topological calculations, as it is essential for capturing the SOC-driven band inversions and avoided crossings responsible for nontrivial topology in tellurene 10.
@@
In contrast, buckled kagome tellurene exhibits nontrivial $\mathbb{Z}_{2}$ topology of the occupied band. The kagome lattice naturally hosts a dense manifold of near-degenerate bands, and buckling enhances SOC-induced hybridization among them. In the absence of SOC, the system lies close to a multi-band crossing regime, while SOC lifts these degeneracies and opens a bulk gap. The resulting Berry curvature is highly localized and unusually large in magnitude near SOC-gapped avoided crossings, reflecting the strong interband mixing inherent to the kagome geometry. Despite the extreme local curvature, time-reversal symmetry enforces a vanishing Chern number, and the nontrivial topology is instead encoded in the $\mathbb{Z}_{2}=1$ invariant. This phase represents the most pronounced manifestation of SOC-driven topology among the structures considered. Buckled square tellurene also realizes a $\mathbb{Z}_{2}=1$-nontrivial phase, but through a comparatively simpler mechanism. Here, SOC gaps symmetry-allowed near-crossings present in the non-SOC band structure, yielding a well-defined insulating manifold with nontrivial Wilson-loop winding. The Berry curvature is more moderate and spatially confined than in the kagome case, indicating fewer competing near-degeneracies. Importantly, the persistence of $\mathbb{Z}_{2}=1$ under moderate strain demonstrates that the square-lattice nontrivial phase is not accidental but instead represents a stable topological regime tunable by lattice deformation.
@@ 5 Acknowledgements
S2 Role of SOC in the band structure
@@
S3 Topological characterization of one-dimensional systems
In one-dimensional crystalline systems, the topological classification of electronic bands differs fundamentally from that of two- and three-dimensional materials. For nonmagnetic systems with spin–orbit coupling that preserve time-reversal symmetry, the electronic structure belongs to symmetry class AII in the Altland–Zirnbauer classification1, 2. In contrast to higher dimensions, this symmetry class does not support a nontrivial $\mathbb{Z}_{2}$ topological insulating phase in one dimension. Consequently, a quantum spin Hall–type invariant cannot be defined for strictly one-dimensional, time-reversal-invariant band insulators. Instead, the relevant bulk quantity in one dimension is the Berry phase accumulated along the Brillouin zone, commonly referred to as the Zak phase3. This quantity is directly related to the electronic polarization and provides the appropriate framework for analyzing the topology of one-dimensional insulating systems.
@@
Within the modern theory of polarization, the Zak phase is related to the bulk electronic polarization $P$ via 4, 5
@@
The Zak phase is, in general, not a topological invariant, as its value depends on the choice of real-space origin and gauge. However, in the presence of spatial inversion symmetry or chiral symmetry, the Zak phase becomes quantized to values of 0 or $\pi$ (mod $2\pi$)6, 7. In such cases, the quantized Zak phase serves as a symmetry-protected topological invariant and enforces the existence of robust boundary phenomena, such as topological end states.

### v3

@@ 3 Results and Discussions
where $\Omega_{xy}(\mathbf{k})$ is the Berry curvature summed over the occupied bands, $\Omega_{xy}(\mathbf{k})=\sum_{n\in\mathrm{occ}}\Omega_{n,xy}(\mathbf{k})$. In our implementation, $\Omega_{n,xy}(\mathbf{k})$ is evaluated using the Kubo (velocity-matrix) expression, as implemented in WannierTools,
@@
The Berry curvature and Chern number were computed using tight-binding Hamiltonians constructed from maximally localized Wannier functions, ensuring an accurate interpolation of the electronic structure on dense $\mathbf{k}$-point meshes. Spin–orbit coupling was included in all topological calculations, as it is essential for capturing the SOC-driven band inversions and avoided crossings responsible for nontrivial topology in tellurene 52.
@@
In contrast, buckled kagome tellurene exhibits nontrivial $\mathbb{Z}_{2}$ topology of the occupied band. The kagome lattice naturally hosts a dense manifold of near-degenerate bands, and buckling enhances SOC-induced hybridization among them. In the absence of SOC, the system lies close to a multi-band crossing regime, while SOC lifts these degeneracies and opens a bulk gap. The resulting Berry curvature is highly localized and unusually large in magnitude near SOC-gapped avoided crossings, reflecting the strong interband mixing inherent to the kagome geometry. Despite the extreme local curvature, time-reversal symmetry enforces a vanishing Chern number, and the nontrivial topology is instead encoded in the $\mathbb{Z}_{2}=1$ invariant. This phase represents the most pronounced manifestation of SOC-driven topology among the structures considered.
Buckled square tellurene also realizes a $\mathbb{Z}_{2}=1$-nontrivial phase, but through a comparatively simpler mechanism. Here, SOC gaps symmetry-allowed near-crossings present in the non-SOC band structure, yielding a well-defined insulating manifold with nontrivial Wilson-loop winding. The Berry curvature is more moderate and spatially confined than in the kagome case, indicating fewer competing near-degeneracies. Importantly, the persistence of $\mathbb{Z}_{2}=1$ under moderate strain demonstrates that the square-lattice nontrivial phase is not accidental but instead represents a stable topological regime tunable by lattice deformation.
@@ 6 Thermal properties
7 Role of SOC in the band structure
@@
8 Topological characterization of one-dimensional systems
In one-dimensional crystalline systems, the topological classification of electronic bands differs fundamentally from that of two- and three-dimensional materials. For nonmagnetic systems with spin–orbit coupling that preserve time-reversal symmetry, the electronic structure belongs to symmetry class AII in the Altland–Zirnbauer classification59, 56. In contrast to higher dimensions, this symmetry class does not support a nontrivial $\mathbb{Z}_{2}$ topological insulating phase in one dimension. Consequently, a quantum spin Hall–type invariant cannot be defined for strictly one-dimensional, time-reversal-invariant band insulators. Instead, the relevant bulk quantity in one dimension is the Berry phase accumulated along the Brillouin zone, commonly referred to as the Zak phase68. This quantity is directly related to the electronic polarization and provides the appropriate framework for analyzing the topology of one-dimensional insulating systems.
@@ 8 Topological characterization of one-dimensional systems
Within the modern theory of polarization, the Zak phase is related to the bulk electronic polarization $P$ via 32, 54
@@
The Zak phase is, in general, not a topological invariant, as its value depends on the choice of real-space origin and gauge. However, in the presence of spatial inversion symmetry or chiral symmetry, the Zak phase becomes quantized to values of 0 or $\pi$ (mod $2\pi$)55, 26. In such cases, the quantized Zak phase serves as a symmetry-protected topological invariant and enforces the existence of robust boundary phenomena, such as topological end states.